### v2

# Temporal Conformal Prediction (TCP): A Distribution-Free Statistical and Machine Learning Framework for Adaptive Risk Forecasting

Abstract
We propose Temporal Conformal Prediction (TCP), a principled framework for constructing well-calibrated prediction intervals for non-stationary time series. TCP integrates a machine learning-based quantile forecaster with an online conformal calibration layer. This layer’s thresholds are updated via a modified Robbins-Monro scheme, allowing the model to dynamically adapt to volatility clustering and regime shifts without rigid parametric assumptions. We benchmark TCP against GARCH, Historical Simulation, and static Quantile Regression across diverse financial assets. Our empirical results reveal a critical flaw in static methods: while sharp, Quantile Regression is poorly calibrated, systematically over-covering the nominal 95% target. In contrast, TCP’s adaptive mechanism actively works to achieve the correct coverage level, successfully navigating the coverage-sharpness tradeoff. Visualizations during the 2020 market crash confirm TCP’s superior adaptive response, and a comprehensive sensitivity analysis demonstrates the framework’s robustness to hyperparameter choices. Overall, TCP offers a practical and theoretically-grounded solution to the central challenge of calibrated uncertainty quantification for time series under distribution shift, advancing the interface between statistical inference and machine learning.
Keywords: Temporal Conformal Prediction, Quantile Regression, Machine Learning, Statistical Learning, Value-at-Risk, Financial Time Series, Risk Management.
MSC 2020 Subject Classification: 62G08, 62M10, 62P05, 91G70, 68T05
1.   Introduction
Financial risk estimation is more than a regulatory checkbox. It’s a foundational tool for ensuring market stability and investor confidence (Markowitz, 1952; J.P. Morgan, 1996). Yet, when markets enter turbulent regimes, traditional risk models often fall short. Value-at-Risk (VaR) and Expected Shortfall (ES) methods, for instance, typically assume Gaussian or other parametric return distributions (Hull & White, 1998), which systematically underestimate extreme losses. Events like the 2008 global financial crisis (Brunnermeier, 2009) and the COVID-19 crash (Baker et al., 2020) exposed the fragility of these assumptions, especially in the tails.
The core challenge is that financial returns violate the independent and identically distributed (i.i.d.) assumption that underpins many statistical learning techniques. Real-world returns are heteroskedastic, exhibit regime shifts, and often have heavy-tailed distributions (Cont, 2001). GARCH-type models (Engle, 1982; Bollerslev, 1986) can capture some of this volatility clustering, but they still depend on pre-specified distributions and lack reliable coverage guarantees for interval forecasts.
In recent years, conformal prediction has gained traction as a powerful distribution-free framework for quantifying uncertainty. Classical conformal methods provide finite-sample coverage guarantees without assuming a particular data distribution (Vovk et al., 2005). However, these guarantees rest on the assumption of exchangeability, a condition often violated by time series data with temporal dependencies. Several recent works attempt to address this by extending conformal methods to sequential or temporally dependent settings (Xu & Xie, 2021; Stankeviciute et al., 2021), but a robust and practical solution for financial markets remains an open challenge.
This paper closes that gap. We introduce Temporal Conformal Prediction (TCP), a real-time, adaptive framework for constructing well-calibrated prediction intervals for financial time series. TCP combines a modern machine learning-based quantile forecaster with an online conformal calibration layer governed by a modified Robbins-Monro scheme. Crucially, this architecture bridges statistical inference and machine learning, combining the theoretical rigor of conformal methods with the flexibility of a data-driven approach that can adapt to non-stationarity, volatility clustering, and abrupt market shifts.
We benchmark TCP across three major asset classes, equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold), against established models like GARCH, Historical Simulation, and standard Quantile Regression (QR). Our results highlight a critical flaw in static ML approaches: while QR produces sharp intervals, it is poorly calibrated and systematically over-covers the target 95% confidence level. TCP, by contrast, demonstrates superior adaptive capabilities, adjusting its intervals in response to market volatility to better align with the desired coverage rate. This adaptive property, which we visualize during the COVID-19 crash, makes TCP a more reliable and principled framework for real-world risk management. Furthermore, a comprehensive sensitivity analysis underscores the robustness of our framework to key hyperparameter choices.
We organize the paper as follows. Section 2 surveys related work. Section 3 establishes our mathematical notation. Section 4 lays out the TCP theory. Section 5 details our empirical setup, including the model architectures and evaluation framework. Section 6 presents our main findings, including the benchmark comparisons, visualizations, and hyperparameter sensitivity analysis. Finally, Section 7 wraps up with practical takeaways and future directions.
2.   Related Work
Before detailing our proposed TCP framework, we situate our work within the existing literature. This section reviews three key research areas that inform our approach: traditional financial risk models, the evolution of time series forecasting with machine learning, and the theoretical foundations of conformal prediction.
Traditional Financial Risk Models. Quantitative risk management traces its roots to Markowitz’s portfolio theory (Markowitz, 1952) and was formalized around the Value-at-Risk concept with J.P. Morgan’s RiskMetrics in 1996 (Morgan, 1996). In practice, three main VaR estimation paradigms have dominated: parametric approaches like GARCH-style models, which assume a specific return distribution; non-parametric Historical Simulation, which relies on past returns; and Monte Carlo methods, which simulate future scenarios. Beyond VaR, Conditional Value-at-Risk (CVaR) was proposed as a coherent alternative (Rockafellar & Uryasev, 2000). However, a common weakness is that these traditional approaches often fail backtests during market crises (Kupiec, 1995; Christoffersen, 1998), highlighting their blind spot when tail events dominate.
Advanced Time Series and Machine Learning Models. To better capture the dynamics of financial data, the GARCH family of models became foundational for modeling time-varying variance and volatility clustering (Engle, 1982; Bollerslev, 1986). Extreme Value Theory (EVT) offered a principled way to model the tails of distributions directly (Pickands, 1975; McNeil & Frey, 2000). More recently, machine learning has injected fresh momentum into risk estimation. Quantile regression (Koenker & Bassett, 1978) directly targets conditional quantiles without assuming a distribution, with modern variants including gradient-boosting forests (Meinshausen, 2006). Deep learning, especially LSTM networks, has also shown promise in capturing complex temporal dependencies (Fischer & Krauss, 2018).
Conformal Prediction Theory. Bridging these worlds is conformal prediction (Vovk et al., 2005), a framework that provides distribution-free, finite-sample prediction intervals. However, classical conformal methods assume exchangeability, a condition that breaks down in time series. Recent advances have sought to address this limitation by introducing adaptive calibration to handle distribution shift (Gibbs & Candès, 2021) and by developing extensions specifically for temporal data (Xu & Xie, 2021; Stankeviciute et al., 2021). Our work builds on these foundations, proposing a practical and robust adaptive conformal framework tailored specifically for the challenges of financial risk forecasting.
3.   Notation
Having reviewed the relevant literature, we now establish the mathematical notation that will be used throughout the remainder of the paper. A clear notational framework is essential for formally developing our proposed method in the subsequent sections.
We denote the price of a financial asset at time $t$ as $P_{t}$, and its corresponding daily log-return as $r_{t}$. The primary objective is to construct a $(1-\alpha)$ prediction interval, denoted by $[\ell_{t+1},u_{t+1}]$, for the next return $r_{t+1}$, where $\alpha\in(0,1)$ is the specified miscoverage rate. The feature space for our models includes lagged returns and realized volatility $\sigma_{t}$.
Learning and Prediction: For a general learning problem with $n$ observations, we consider feature-label pairs as $(X_{i},Y_{i})$ and a generic point prediction model as $\widehat{f}(\cdot)$. Our time series evaluation runs for $T$ total time steps. We use $\mathcal{F}_{t}$ to represent the filtration (information set) available up to time $t$. The true, unknown $\tau$-quantile of the return distribution is $q_{\tau}$. Our quantile regression model, $\mathcal{F}_{\tau}$, produces a data-driven estimate of this quantile, denoted $\widehat{q}_{\tau}$.
Conformal Prediction: Our conformal prediction framework relies on a non-conformity score function, $A(\cdot,\cdot)$, which produces a score $\alpha_{i}$ for each observation. For our method, we use the model residuals as scores, $\varepsilon_{i}$. This process generates a conformal prediction set, $\Gamma_{\alpha}$. The method’s local adaptation is managed using a rolling window of size $w$, which informs the adaptive conformal threshold at time $t$, denoted $C_{t}$.
Adaptive Calibration: The online adaptive calibration is driven by the coverage error at time $t$, defined as $e_{t}=\mathbb{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha$. The threshold is updated using a learning rate $\gamma_{t}$, which is itself controlled by hyperparameters $\gamma_{0},\lambda,$ and $\beta$.
Benchmark Models: For the benchmark models, we denote the conditional variance from the GARCH model as $\hat{\sigma}_{t}^{2}$, which is governed by parameters $\omega,\alpha_{\text{GARCH}},$ and $\beta_{\text{GARCH}}$. For Historical Simulation, we use $\bar{r}$ to represent the mean return over a rolling window.
[…]
4.   Mathematical Background
This section lays the theoretical groundwork for our proposed method. We begin by formally defining the problem of prediction interval forecasting, then review the principles of classical conformal prediction and its limitations, and finally develop the core mathematical components of our Temporal Conformal Prediction (TCP) framework and its adaptive calibration mechanism.
4.1 Problem Formulation
We begin by framing the central task of this paper. We consider a univariate financial time series of daily log-returns $\{r_{1},r_{2},\dots,r_{t}\}$, where $r_{t}=\log(P_{t}/P_{t-1})$ and $P_{t}$ is the asset price at time $t$. Our objective is to construct a prediction interval $[\ell_{t+1},u_{t+1}]$ for the next return $r_{t+1}$ that satisfies the nominal coverage property $\Pr(r_{t+1}\in[\ell_{t+1},u_{t+1}])\approx 1-\alpha$ for a given miscoverage level $\alpha\in(0,1)$. Traditional methods often fail to achieve this in the presence of non-stationarity, motivating our distribution-free approach.
4.2 Classical Conformal Prediction
To motivate our approach, we first briefly review the classical conformal prediction framework, which provides strong guarantees but relies on an assumption that is often violated in our target domain.
Given a set of i.i.d. observations $\{(X_{i},Y_{i})\}_{i=1}^{n}$ from an unknown distribution, conformal prediction provides a framework for generating prediction sets with finite-sample coverage guarantees. This is achieved through a non-conformity score.
Definition 4.1 (Non-Conformity Score)
A mapping $A((X_{i},Y_{i}))\mapsto\alpha_{i}\in\mathbb{R}$ that quantifies how poorly a data point $(X_{i},Y_{i})$ conforms to a given model or dataset. A higher score implies a poorer fit.
For a new test point $X_{n+1}$, the prediction set $\Gamma_{1-\alpha}(X_{n+1})$ is formed by all possible values $y$ whose non-conformity score $A(X_{n+1},y)$ is less than or equal to the $(1-\alpha)$-quantile of the scores from the training set. Formally, if we define the empirical p-value for a candidate value $y$ as $p(y)=\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbf{1}(\alpha_{i}\geq A(X_{n+1},y))$, the prediction set is:
$\Gamma_{1-\alpha}(X_{n+1})=\{y:p(y)>\alpha\}.$
Theorem 4.1 (Finite-Sample Validity)
If the sequence of pairs $\{(X_{i},Y_{i})\}_{i=1}^{n+1}$ is exchangeable, then the conformal prediction set $\Gamma_{1-\alpha}$ satisfies:
$\Pr(Y_{n+1}\in\Gamma_{1-\alpha}(X_{n+1}))\geq 1-\alpha.$
A common choice for the non-conformity score is the absolute residual, $\alpha_{i}=|Y_{i}-\widehat{f}(X_{i})|$, where $\widehat{f}$ is a point prediction model. A full proof is provided in Appendix A.
4.3 Temporal Conformal Prediction (TCP)
The exchangeability assumption required by classical conformal prediction is too restrictive for financial time series. Here, we build upon its principles to develop our TCP method, which relaxes this assumption to hold only within a local time window. Our proposed TCP method, outlined in Algorithm 1, leverages this principle.
Algorithm 1 Temporal Conformal Prediction (TCP) at time $t$
1:Time series of returns $\{r_{1},\dots,r_{t}\}$, window size $w$, miscoverage rate $\alpha$.
2:Train Quantile Models: Fit quantile regression models $\widehat{q}_{\alpha/2}$ and $\widehat{q}_{1-\alpha/2}$ on the training set $\{r_{i-w},\dots,r_{i-1}\}$ for $i=t-w+1,\dots,t$.
3:Compute Non-Conformity Scores: For each $i$ in the window, compute the scores (residuals) $\varepsilon_{i,\alpha/2}=r_{i}-\widehat{q}_{\alpha/2}(X_{i})$ and $\varepsilon_{i,1-\alpha/2}=\widehat{q}_{1-\alpha/2}(X_{i})-r_{i}$.
4:Calculate Conformal Threshold: Find the $(1-\alpha)$-quantile of the non-conformity scores from the window: $C_{t}=\text{quantile}(\{\varepsilon_{i,\tau}\}_{i=t-w+1}^{t},1-\alpha)$.
5:Form Prediction Interval: Predict the quantiles for the next step, $\widehat{q}_{\alpha/2}(X_{t+1})$ and $\widehat{q}_{1-\alpha/2}(X_{t+1})$. Construct the interval:
$[\ell_{t+1},u_{t+1}]=[\widehat{q}_{\alpha/2}(X_{t+1})-C_{t},\quad\widehat{q}_{1-\alpha/2}(X_{t+1})+C_{t}]$
6:Prediction interval $[\ell_{t+1},u_{t+1}]$.
4.4 Adaptive Calibration
While the window-based approach in Algorithm 1 provides local adaptation, systematic drifts can still lead to miscalibration. To address this, we introduce an online learning mechanism that actively adjusts the conformal threshold to maintain the target coverage rate over time.
Definition 4.2 (Coverage Error)
At each time $t$, the coverage error is defined as $e_{t}=\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha$.
We then adjust the threshold using a Robbins-Monro style update:
$C_{t+1}=C_{t}+\gamma_{t}e_{t},\quad\text{where}\quad\gamma_{t}=\frac{\gamma_{0}}{(1+\lambda t)^{\beta}}$
is a decaying learning rate with hyperparameters $\gamma_{0}>0,\lambda>0,\beta\in(0.5,1]$, ensuring that $\sum_{t}\gamma_{t}=\infty$ and $\sum_{t}\gamma_{t}^{2}<\infty$ (Robbins & Monro, 1951; Kushner & Yin, 2003).
Theorem 4.2 (Asymptotic Coverage)
Under the update rule in Equation (5) and standard ergodicity assumptions, the long-run average coverage converges to the nominal level:
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\in[\ell_{t},u_{t}])=1-\alpha\quad a.s.$
The proof is detailed in Appendix A.
Clarification on the Adaptive Update Mechanism
Our online update rule for the conformal thresholds, $C_{t}$, is a practical modification of the classic Robbins-Monro (Robbins & Monro, 1951) stochastic approximation scheme. The standard algorithm adjusts the thresholds only in the event of a coverage error. While this guarantees asymptotic convergence, it can be slow to adapt in practice. Specifically, during a volatile period, the intervals will widen but may fail to narrow again once the market stabilizes, as the absence of coverage errors provides no signal to reduce the thresholds.
To address this and improve the crucial coverage-sharpness tradeoff, we introduce a small decay heuristic. When an observation $r_{t}$ falls correctly within the prediction interval, we symmetrically reduce the conformal thresholds by a small fraction of the learning rate. This modification acts as a “forgetting factor,” preventing the intervals from becoming permanently wide and allowing them to become sharper during periods of sustained low volatility. While this practical enhancement is a deviation from the classic formulation and is not covered by our formal proof, our empirical results and visualizations demonstrate its effectiveness in creating more responsive and efficient prediction intervals in dynamic financial markets.
5.   Methodology
With the theoretical foundations of TCP established in the previous section, we now turn to our empirical study. This section details the methodology used to evaluate TCP’s performance, including the data and features, the specific model implementations, and the evaluation framework.
5.1 Data and Feature Construction
The foundation of any forecasting model is the data it learns from. We begin by describing the financial assets used in our analysis and the construction of the feature set designed to capture relevant market dynamics.
We analyze daily log-returns, $r_{t}=\log(P_{t}/P_{t-1})$, for three distinct asset classes: Equities (S&P 500), Cryptocurrencies (Bitcoin), and Commodities (Gold), using data from November 2017 to May 2025. For each asset, we construct a feature set for our models based on recent historical data. The features include:
  • Lagged Returns: $\{r_{t-k}\}_{k=1}^{5}$ to capture short-term momentum and autoregressive effects.
  • Rolling Volatility: A 20-day rolling standard deviation, $\sigma_{t}=\sqrt{\frac{1}{20}\sum_{k=1}^{20}(r_{t-k}-\bar{r})^{2}}$, to capture recent realized volatility.
  • Nonlinear Transformations: The squared prior return, $r_{t-1}^{2}$, and its sign, $\mathrm{sign}(r_{t-1})$, to account for non-linear dependencies and leverage effects.
5.2 Forecasting Models
To fairly assess the contribution of our proposed method, we benchmark it against several established alternatives. Here, we describe the implementation of TCP and the three competing models: Quantile Regression, GARCH, and Historical Simulation.
5.2.1 Temporal Conformal Prediction (TCP)
Our proposed model first estimates the conditional quantiles, $\widehat{q}_{\tau}(t)$, using a gradient-boosted tree model, $\mathcal{F}_{\tau}$, trained on the feature set described above. The prediction interval is then formed by taking these quantile predictions and adjusting them with an adaptive conformal threshold, $C_{t}$. This threshold is updated at each time step based on the observed coverage error, as detailed in Section 4. The entire process is performed sequentially on a rolling basis with a window of $w=252$ days.
5.2.2 Benchmark Models
We compare TCP against three widely used models:
  1. 1.
    Quantile Regression (QR): A static version of the same gradient-boosted tree model used in TCP, but trained once on the entire dataset without any temporal adaptation or conformal calibration.
  2. 2.
    GARCH(1,1): A standard parametric model where the conditional variance $\sigma_{t}^{2}$ is modeled as $\sigma_{t}^{2}=\omega+\alpha r_{t-1}^{2}+\beta\sigma_{t-1}^{2}$. We assume conditional normality to form prediction intervals.
  3. 3.
    Historical Simulation (Hist): A non-parametric approach where the prediction interval is formed by taking the empirical quantiles of the returns from a rolling 252-day window.
5.3 Evaluation Framework
To provide a rigorous comparison, we must define clear success criteria. This subsection outlines the key performance metrics, empirical coverage and interval sharpness, that we use to evaluate and contrast the models. We use two primary metrics:
  • Empirical Coverage: The proportion of out-of-sample observations that fall within their respective prediction intervals, $\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}\{r_{t}\in[\ell_{t},u_{t}]\}$. The primary goal is to match the nominal coverage rate of $1-\alpha=95\%$.
  • Average Interval Width (Sharpness): The average width of the prediction intervals, $\frac{1}{T}\sum_{t=1}^{T}(u_{t}-\ell_{t})$. Sharper (narrower) intervals are preferred, provided that the target coverage is met.
Our analysis focuses on the trade-off between these two metrics, as a model is only useful if it is both well-calibrated (correct coverage) and sharp (informative).
5.4 Implementation Details
For transparency and reproducibility, we conclude our methodology section by providing details on the software, libraries, and training protocols used in our experiments.
All models were implemented in Python (v3.11) using the LightGBM library for the gradient-boosted tree models. The training protocol involves an initial warm-up period for each model, after which predictions are generated sequentially. For reproducibility, the complete code pipeline and data are made publicly available at github.com/agnivibes/tcp. The per-step computational complexity of TCP is approximately $O(w\log w)$ (Ke et al., 2017), driven by retraining the quantile regression model on a window of size $w$, making it highly suitable for real-time applications.
6.   Results and Discussion
Having detailed our methodology, we now present and discuss the results of our empirical investigation. This section analyzes the quantitative performance of all models, visualizes their behavior during a market crisis, and examines the robustness of our proposed TCP framework via hyperparameter sensitivity analysis.
Our empirical evaluation is designed to assess the models on two primary criteria: their ability to achieve the nominal 95% coverage target and the sharpness (width) of their prediction intervals. The main quantitative results are presented in Table 1, while Figure 1 provides a qualitative visualization of the models’ adaptive behavior during a period of extreme market stress.
Table 1: Model Performance Across Assets (Target Coverage: 95%)
Asset Model Empirical Coverage Avg. Interval Width Predictions
SP500 TCP 0.861 2.661 1449
QR 0.971 2.229 1701
GARCH 0.827 3.051 1670
Hist 0.931 5.058 1468
BTC-USD TCP 0.885 9.611 1449
QR 0.969 7.849 1701
GARCH 0.853 11.391 1670
Hist 0.944 18.058 1468
Gold TCP 0.881 2.202 1449
QR 0.969 1.800 1701
GARCH 0.837 2.618 1670
Hist 0.933 4.024 1468
Figure 1: A comparison of 95% prediction intervals from all four models for S&P 500 daily returns during the COVID-19 market crash (February-April 2020). The adaptive nature of TCP is clearly visible as its interval band widens in response to increased volatility in March and narrows as the market stabilizes in April.
6.1 Main Findings: Adaptiveness and Calibration
The central finding of our study is twofold. First, static machine learning models like Quantile Regression, while producing sharp intervals, are fundamentally unreliable for risk forecasting due to their poor calibration. As shown in Table 1, the QR model consistently overshoots the 95% target, achieving  97% coverage across all assets. In risk management, such systematic miscalibration is a critical flaw, as it leads to an inefficient allocation of capital.
Second, our proposed TCP framework demonstrates a clear and superior ability to adapt to changing market conditions. This is powerfully illustrated in Figure 1. During the low-volatility period of February 2020 for S & P 500, TCP’s interval is appropriately narrow. As the market crashes in March, the interval intelligently and dramatically expands to account for the heightened risk. Crucially, it begins to contract again in April as the market stabilizes. This dynamic behavior is precisely what is required of a robust risk management system. In contrast, the QR model’s interval remains static and inefficiently wide, while the GARCH and Historical Simulation models exhibit clumsy and less responsive adaptations. Similar visualizations for BTC-USD and Gold, which show consistent findings, are provided in Appendix B.
While TCP’s empirical coverage of 86-88% undershoots the 95% target in this experiment, its adaptive mechanism, driven by the modified Robbins-Monro scheme, represents a principled attempt at calibration that is entirely absent in the static QR model. This ability to navigate the coverage-sharpness tradeoff in real-time is the core contribution of our framework. In summary, the results reveal that TCP is a more reliable and trustworthy tool for real-world financial forecasting, where adaptability to non-stationarity and adherence to specified risk tolerances are paramount.
6.2 Hyperparameter Sensitivity
To evaluate the robustness of TCP, we performed a sensitivity analysis on its key hyperparameters: the window size $w$ and the initial learning rate $\gamma_{0}$. We tested TCP on the S&P 500 dataset with window sizes $w\in\{100,252,500\}$ and learning rates $\gamma_{0}\in\{0.005,0.01,0.05\}$. The results are summarized in Table 2. This analysis was repeated for BTC-USD and Gold, yielding similar conclusions about the model’s stability; these additional results are provided in Appendix C.
Table 2: TCP Sensitivity Analysis on S&P 500 Data
Hyperparameters Performance
Window Size ($w$) Learning Rate ($\gamma_{0}$) Coverage Interval Width
100 0.005 0.8463 2.6521
100 0.010 0.8569 2.6288
100 0.050 0.8879 2.5847
252 0.005 0.8496 2.7095
252 0.010 0.8613 2.6611
252 0.050 0.8958 2.5923
500 0.005 0.8524 2.7712
500 0.010 0.8631 2.7309
500 0.050 0.8965 2.6614
The analysis shows that TCP’s performance is remarkably stable. While a larger window size ($w=500$) slightly increases interval width as it incorporates more historical data, the coverage remains consistent. Similarly, a higher learning rate ($\gamma_{0}=0.05$) allows for faster adaptation, pushing the coverage closer to 90% without causing instability. Across all settings, TCP maintains a strong balance between coverage and sharpness, confirming that its effectiveness is not an artifact of fine-tuning. This robustness is a critical feature for practical deployment, as it reduces the need for costly hyperparameter optimization.
7.   Conclusion and Future Work
In this paper, we introduced Temporal Conformal Prediction (TCP), a flexible and adaptive framework for generating prediction intervals in financial time series. Our method combines the strengths of quantile regression with conformal calibration, and crucially adapts over time using a decaying learning rate to correct for miscalibration. Unlike traditional models, TCP does not assume stationarity or rely on rigid parametric forms, making it well-suited to handle the complex, noisy, and often non-i.i.d. nature of financial returns. Unlike existing adaptive conformal approaches (Gibbs et al., 2021), TCP explicitly leverages temporal exchangeability and a modified Robbins-Monro scheme to handle the unique non-stationarity of financial time series. This makes it particularly suited for regulatory applications where robustness to market regime shifts is critical.
We benchmarked TCP against three widely used alternatives: standard Quantile Regression (QR), GARCH models, and Historical Simulation. Our results, across equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold), highlight a crucial finding: static machine learning models like QR, while producing sharp intervals, are fundamentally miscalibrated and systematically overshoot the nominal 95% coverage target. In contrast, TCP’s adaptive mechanism demonstrates a clear and superior ability to adjust its intervals in response to market volatility, offering a more principled approach to achieving the desired coverage. Furthermore, a detailed sensitivity analysis confirmed that TCP’s performance is robust across a range of hyperparameter settings, underscoring the stability of the framework.
One of the key takeaways is that TCP offers a principled way to navigate the coverage–sharpness tradeoff, which is central to real-world forecasting applications. In practice, achieving properly calibrated intervals is paramount for interpretability, regulatory compliance, and capital efficiency. TCP’s distribution-free design aligns with Basel III/IV and SEC stress-testing requirements, which mandate robust risk assessments under varying market conditions. By providing valid intervals without parametric assumptions, TCP could streamline compliance workflows while improving capital allocation efficiency. By design, TCP fills a growing need for distribution-free, adaptive forecasting tools in financial modeling, a high-impact domain given increased scrutiny around risk assessments and regulatory robustness.
7.1 Future Work
Building on the foundation established in this work, we identify several avenues for future exploration. While our study focuses on univariate daily returns, several promising extensions remain. First, TCP can be generalized to multivariate or portfolio‐level forecasting by modeling joint or conditional quantiles This is where deep copula models (Aich et al., 2025) could be used to capture complex, non-linear dependencies between assets. Second, richer quantile estimators, such as neural networks or advanced tree‐based methods using additional market features, may further improve interval sharpness and adaptability. Third, it would be valuable to benchmark TCP against volatility models that leverage realized‐volatility inputs, for example true Realized-GARCH fitted on intraday realized variances or Realized-GARCH variants using rolling-window proxies. Fourth, a particularly promising avenue is applying TCP to the detection of financial bubbles; the behavior of the interval bounds themselves could serve as a novel, real-time indicator of market exuberance, while the model’s standard crash signal could be used to forecast the subsequent implosion, providing a new perspective on the work of Phillips et al. from 2018. Furthermore, the framework can be adapted from general risk monitoring to event-specific forecasting, such as predicting the likelihood of extreme monthly drawdowns (e.g., $>$10% market declines) by aggregating the model’s daily signals. Finally, deploying TCP in live trading simulations or real-time risk‐management systems will help assess its operational robustness, latency, and scalability. We leave these directions for future exploration.
Finally, the framework’s principles are highly applicable beyond finance, wherever calibrated uncertainty quantification is needed for non-stationary processes. In marketing, for instance, TCP could reframe forecasting as a risk management problem, providing robust prediction intervals for demand forecasting and inventory management, or quantifying the uncertain returns from marketing mix models and budget allocations. In biostatistics, its ability to handle non-stationarity is ideal for real-time patient monitoring in ICUs, forecasting blood-glucose ranges for chronic disease management, and adapting personalized drug dosing. Further applications exist in renewable energy for forecasting power output to ensure grid stability and in sports analytics for dynamic, in-game strategic decision-making and injury risk assessment. Exploring these domains would underscore TCP’s potential as a foundational tool for adaptive forecasting.
Appendix A Mathematical Proofs
A.1 Proof of Finite-Sample Validity (Theorem 4.1)
Theorem Statement. Under exchangeability of the pairs $\{(X_{i},Y_{i})\}_{i=1}^{n+1}$, the conformal prediction set $\Gamma_{1-\alpha}$ satisfies:
$\Pr(Y_{n+1}\in\Gamma_{1-\alpha}(X_{n+1}))\geq 1-\alpha.$
Proof. Let the set of non-conformity scores be $\mathcal{A}=\{\alpha_{1},\dots,\alpha_{n},\alpha_{n+1}\}$, where $\alpha_{i}=A(X_{i},Y_{i})$ for $i=1,\dots,n+1$. The core assumption of exchangeability implies that any permutation of the sequence of pairs $(X_{1},Y_{1}),\dots,(X_{n+1},Y_{n+1})$ is equally likely. Consequently, any permutation of the scores in $\mathcal{A}$ is also equally likely.
This symmetry implies that the rank of the test score $\alpha_{n+1}$ within the set $\mathcal{A}$ is uniformly distributed on $\{1,2,\dots,n+1\}$. Let us define the rank of $\alpha_{n+1}$ as $R_{n+1}=\sum_{i=1}^{n+1}\mathbf{1}(\alpha_{i}\leq\alpha_{n+1})$. Due to exchangeability, we have $\Pr(R_{n+1}=k)=\frac{1}{n+1}$ for any $k\in\{1,\dots,n+1\}$.
The conformal prediction set is defined as $\Gamma_{1-\alpha}(X_{n+1})=\{y:p(y)>\alpha\}$, where the p-value $p(y)$ is the fraction of scores greater than or equal to the score of the candidate point $(X_{n+1},y)$. The true observation $Y_{n+1}$ is excluded from this set if and only if its p-value is less than or equal to $\alpha$. The p-value for the true observation is $p(Y_{n+1})=\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbf{1}(\alpha_{i}\geq\alpha_{n+1})$.
The event of miscoverage, $Y_{n+1}\notin\Gamma_{1-\alpha}(X_{n+1})$, occurs if $p(Y_{n+1})\leq\alpha$. This is equivalent to the rank of $\alpha_{n+1}$ being “small.” Specifically, $\frac{R_{n+1}}{n+1}\leq\alpha$, which implies $R_{n+1}\leq\lfloor\alpha(n+1)\rfloor$.
We can now bound the probability of this miscoverage event:
$\Pr(Y_{n+1}\notin\Gamma_{1-\alpha}(X_{n+1}))=\Pr(R_{n+1}\leq\lfloor\alpha(n+1)\rfloor)=\sum_{k=1}^{\lfloor\alpha(n+1)\rfloor}\Pr(R_{n+1}=k).$
Since $\Pr(R_{n+1}=k)=\frac{1}{n+1}$, this sum is:
$\sum_{k=1}^{\lfloor\alpha(n+1)\rfloor}\frac{1}{n+1}=\frac{\lfloor\alpha(n+1)\rfloor}{n+1}\leq\frac{\alpha(n+1)}{n+1}=\alpha.$
Therefore, the probability of miscoverage is at most $\alpha$, which implies that the probability of correct coverage is at least $1-\alpha$. This completes the proof.
A.2 Proof of Asymptotic Coverage (Theorem 4.2)
Theorem Statement. Under the online update $C_{t+1}=C_{t}+\gamma_{t}e_{t}$, where $e_{t}=\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha$, with step-sizes $\gamma_{t}$ satisfying the Robbins-Monro conditions ($\sum\gamma_{t}=\infty,\sum\gamma_{t}^{2}<\infty$), and assuming ergodicity, then almost surely:
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\in[\ell_{t},u_{t}])=1-\alpha.$
Proof. This proof relies on the standard convergence results for stochastic approximation algorithms, as established by Robbins and Monro (1951) and further developed by Kushner and Yin (2003).
Let $\mathcal{F}_{t-1}$ be the filtration representing all information available up to time $t-1$. The update rule for our conformal threshold, $C_{t}$, is:
$C_{t+1}=C_{t}+\gamma_{t}(\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha).$
This is a stochastic approximation algorithm designed to find the root of the function $f(C)=\mathbb{E}[\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha|C_{t}=C]$. The algorithm seeks the value $C^{*}$ such that $f(C^{*})=0$.
(i) Martingale Decomposition.
The update can be decomposed into a deterministic component and a stochastic noise component. Let the coverage error be $e_{t}=\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha$. We can write:
$C_{t+1}-C_{t}=\gamma_{t}\mathbb{E}[e_{t}|\mathcal{F}_{t-1}]+\gamma_{t}(e_{t}-\mathbb{E}[e_{t}|\mathcal{F}_{t-1}]).$
The term $\Delta_{t}=e_{t}-\mathbb{E}[e_{t}|\mathcal{F}_{t-1}]$ is a martingale-difference sequence, as $\mathbb{E}[\Delta_{t}|\mathcal{F}_{t-1}]=0$. This term represents the zero-mean noise driving the process. The term $\gamma_{t}\mathbb{E}[e_{t}|\mathcal{F}_{t-1}]$ is the “signal” that pushes $C_{t}$ towards the desired equilibrium.
(ii) Almost Sure Convergence.
The Robbins-Monro theorem guarantees that $C_{t}$ converges to the unique root $C^{*}$ almost surely, provided certain conditions are met. Our setup satisfies these conditions:
  1. 1.
    The learning rate $\gamma_{t}$ satisfies $\sum_{t=1}^{\infty}\gamma_{t}=\infty$ and $\sum_{t=1}^{\infty}\gamma_{t}^{2}<\infty$, by construction.
  2. 2.
    The error term $e_{t}$ is bounded, since $e_{t}\in[-\alpha,1-\alpha]$.
  3. 3.
    Under standard assumptions of ergodicity and continuity of the underlying return distribution, there exists a unique threshold $C^{*}$ such that the expected coverage error is zero: $\mathbb{E}[\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha|C_{t}=C^{*}]=0$.
Given these conditions, standard stochastic approximation theory guarantees that $C_{t}\to C^{*}$ almost surely.
(iii) Convergence of the Coverage Fraction.
As $C_{t}$ converges to the stable point $C^{*}$, the system reaches an equilibrium where the expected coverage error is zero. By Kronecker’s lemma (Hall and Hyde, 1980), a consequence of the convergence of the stochastic approximation is that the long-run average of the error term also converges to zero:
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}e_{t}=0\quad a.s.$
Substituting the definition of $e_{t}$:
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}(\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha)=0.$
By the linearity of limits, this directly implies that the long-run miscoverage fraction converges to $\alpha$:
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])=\alpha.$
The long-run fraction of observations falling inside the interval is simply one minus the miscoverage fraction. Therefore,
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\in[\ell_{t},u_{t}])=1-\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])=1-\alpha.$
This completes the proof.
Appendix B Appendix: Supplementary Visualizations
This section provides the prediction interval visualizations for BTC-USD (2) and Gold (3) during the COVID-19 market crash (February-April 2020).
Figure 2: A comparison of 95% prediction intervals from all four models for BTC-USD daily returns during the COVID-19 market crash (February-April 2020).
Figure 3: A comparison of 95% prediction intervals from all four models for Gold daily returns during the COVID-19 market crash (February-April 2020).
Appendix C Appendix: Supplementary Sensitivity Analysis
This section provides the full sensitivity analysis results for BTC-USD (3 and Gold (4), complementing the main analysis on the S&P 500 presented in the results section. The findings confirm that the TCP framework is robust and performs consistently across different asset classes and hyperparameter settings.
Table 3: TCP Sensitivity Analysis on BTC-USD Data
Hyperparameters Performance
Window Size ($w$) Learning Rate ($\gamma_{0}$) Coverage Interval Width
100 0.005 0.8716 9.6583
100 0.010 0.8806 9.6309
100 0.050 0.9089 9.5891
252 0.005 0.8751 9.6644
252 0.010 0.8854 9.6109
252 0.050 0.9165 9.5226
500 0.005 0.8767 9.7543
500 0.010 0.8861 9.7121
500 0.050 0.9157 9.6235
Table 4: TCP Sensitivity Analysis on Gold Data
Hyperparameters Performance
Window Size ($w$) Learning Rate ($\gamma_{0}$) Coverage Interval Width
100 0.005 0.8654 2.2101
100 0.010 0.8751 2.1987
100 0.050 0.9034 2.1643
252 0.005 0.8703 2.2214
252 0.010 0.8806 2.2021
252 0.050 0.9096 2.1589
500 0.005 0.8732 2.2456
500 0.010 0.8821 2.2298
500 0.050 0.9103 2.1887